OpenStudy (anonymous):

(3x+3)/(x^2+2x+1)+(x-1)/(x^2-1) help please..... I think that the denomanators would be (x+1)(x+1) right? and (x-1) (x-1)

OpenStudy (anonymous):

\[\dfrac{(3x+3)}{(x^2+2x+1)}+\dfrac{(x-1)}{(x^2-1)}\] You're right about the denominator of the first fraction, that does factor to \((x+1)(x+1)\). The second one, though, doesn't factor to \((x-1)(x-1)\), though, because that would be \((x^2-2x+1)\). It does factor to \((x+1)(x-1)\), from the difference of squares formula. That means you're left with: \[\dfrac{(3x+3)}{(x+1)(x+1)}+\dfrac{(x-1)}{(x+1)(x-1)}\] Can you solve it from there?

OpenStudy (anonymous):

would the (x-1) on top cancell out with the (x-1) on the bottom?

OpenStudy (anonymous):

It would. The next step after that is to make both denominators the same so that you can add the fractions together.

OpenStudy (anonymous):

so then would it be 3x+3/x+1

OpenStudy (thomas5267):

HINTS: \[ \frac{(3x+3)}{(x+1)(x-1)}=\frac{3(x+1)}{(x+1)(x+1)} \]

OpenStudy (anonymous):

How did you get there?

OpenStudy (anonymous):

@thomas5267 where did you get the 3 on the top part of the other equation?

OpenStudy (thomas5267):

I factored it out. See this. http://www.purplemath.com/modules/simpfact.htm

OpenStudy (anonymous):

oh okay thanks